\def\cfudot#1{\ifmmode\setbox7\hbox{$\accent"5E#1$}\else
  \setbox7\hbox{\accent"5E#1}\penalty 10000\relax\fi\raise 1\ht7
  \hbox{\raise.1ex\hbox to 1\wd7{\hss.\hss}}\penalty 10000 \hskip-1\wd7\penalty
  10000\box7}
\begin{thebibliography}{CHKV06}

\bibitem[AL08]{AberbachLiAsymptoticVanishingRegular}
Ian~M. Aberbach and Jinjia Li.
\newblock Asymptotic vanishing conditions which force regularity in local rings
  of prime characteristic.
\newblock {\em Math. Res. Lett.}, 15(4):815--820, 2008.

\bibitem[Bha12]{BhattDerivedsplintersinpositivecharacteristic}
Bhargav Bhatt.
\newblock Derived splinters in positive characteristic.
\newblock {\em Compos. Math.}, 148(6):1757--1786, 2012.

\bibitem[BIM19]{BhattIyengarMaRegularRingsPerfectoidAlgebras}
Bhargav Bhatt, Srikanth~B. Iyengar, and Linquan Ma.
\newblock Regular rings and perfect(oid) algebras.
\newblock {\em Communications in Algebra}, 47(6):2367--2383, 2019.

\bibitem[BST15]{BlickleSchwedeTuckerTestAlterations}
Manuel Blickle, Karl Schwede, and Kevin Tucker.
\newblock {$F$}-singularities via alterations.
\newblock {\em Amer. J. Math.}, 137(1):61--109, 2015.

\bibitem[CHKV06]{CorsoHunekeKatzVasconcelos}
Alberto Corso, Craig Huneke, Daniel Katz, and Wolmer~V. Vasconcelos.
\newblock Integral closure of ideals and annihilators of homology.
\newblock In {\em Commutative algebra}, volume 244 of {\em Lect. Notes Pure
  Appl. Math.}, pages 33--48. Chapman \& Hall/CRC, Boca Raton, FL, 2006.

\bibitem[FI03]{FoxbyIyengarDepthAmplitude}
Hans-Bj\o{}rn Foxby and Srikanth Iyengar.
\newblock Depth and amplitude for unbounded complexes.
\newblock In {\em Commutative algebra ({G}renoble/{L}yon, 2001)}, volume 331 of
  {\em Contemp. Math.}, pages 119--137. Amer. Math. Soc., Providence, RI, 2003.

\bibitem[GR70]{GRVanishing}
Hans Grauert and Oswald Riemenschneider.
\newblock Verschwindungss\"atze f\"ur analytische {K}ohomologiegruppen auf
  komplexen {R}\"aumen.
\newblock {\em Invent. Math.}, 11:263--292, 1970.

\bibitem[Hoc75]{HochsterTopicsInTheHomologicalTheory}
Melvin Hochster.
\newblock {\em Topics in the homological theory of modules over commutative
  rings}.
\newblock Published for the Conference Board of the Mathematical Sciences by
  the American Mathematical Society, Providence, R.I., 1975.
\newblock Expository lectures from the CBMS Regional Conference held at the
  University of Nebraska, Lincoln, Neb., June 24--28, 1974, Conference Board of
  the Mathematical Sciences Regional Conference Series in Mathematics, No. 24.

\bibitem[Kov00]{KovacsRat}
S{\'a}ndor~J. Kov{\'a}cs.
\newblock A characterization of rational singularities.
\newblock {\em Duke Math. J.}, 102(2):187--191, 2000.

\bibitem[Kun69]{KunzCharacterizationsOfRegularLocalRings}
Ernst Kunz.
\newblock Characterizations of regular local rings for characteristic {$p$}.
\newblock {\em Amer. J. Math.}, 91:772--784, 1969.

\bibitem[Laz04]{LazarsfeldPositivity2}
Robert Lazarsfeld.
\newblock {\em Positivity in algebraic geometry. {II}}, volume~49 of {\em
  Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of
  Modern Surveys in Mathematics [Results in Mathematics and Related Areas. 3rd
  Series. A Series of Modern Surveys in Mathematics]}.
\newblock Springer-Verlag, Berlin, 2004.
\newblock Positivity for vector bundles, and multiplier ideals.

\bibitem[Lip78]{LipmanDesingularizationOf2Dimensional}
Joseph Lipman.
\newblock Desingularization of two-dimensional schemes.
\newblock {\em Ann. Math. (2)}, 107(1):151--207, 1978.

\bibitem[Mat89]{MatsumuraCommutativeRingTheory}
Hideyuki Matsumura.
\newblock {\em Commutative ring theory}, volume~8 of {\em Cambridge Studies in
  Advanced Mathematics}.
\newblock Cambridge University Press, Cambridge, second edition, 1989.
\newblock Translated from the Japanese by M. Reid.

\bibitem[PS73]{PeskineSzpiroDimensionProjective}
C.~Peskine and L.~Szpiro.
\newblock Dimension projective finie et cohomologie locale. {A}pplications \`a
  la d\'emonstration de conjectures de {M}. {A}uslander, {H}. {B}ass et {A}.
  {G}rothendieck.
\newblock {\em Inst. Hautes \'Etudes Sci. Publ. Math.}, (42):47--119, 1973.

\bibitem[Rob80]{RobertsCohenMacaulayComplexes}
Paul Roberts.
\newblock Cohen-{M}acaulay complexes and an analytic proof of the new
  intersection conjecture.
\newblock {\em J. Algebra}, 66(1):220--225, 1980.

\bibitem[Rob87]{RobertsIntersectionTheorems}
Paul Roberts.
\newblock Le th\'{e}or\`eme d'intersection.
\newblock {\em C. R. Acad. Sci. Paris S\'{e}r. I Math.}, 304(7):177--180, 1987.

\bibitem[Rod88]{RodicioOnaResultAvramov}
Antonio~G. Rodicio.
\newblock On a result of {A}vramov.
\newblock {\em Manuscripta Math.}, 62(2):181--185, 1988.

\bibitem[Sha72]{SharpFinitelyGeneratedmodulesFiniteInjectiveDimension}
Rodney~Y. Sharp.
\newblock Finitely generated modules of finite injective dimension over certain
  {C}ohen-{M}acaulay rings.
\newblock {\em Proc. London Math. Soc. (3)}, 25:303--328, 1972.

\bibitem[ST08]{SchwedeTakagiRationalPairs}
Karl Schwede and Shunsuke Takagi.
\newblock Rational singularities associated to pairs.
\newblock {\em Michigan Math. J.}, 57:625--658, 2008.

\bibitem[Tri94]{TrivediLocalBertini}
Vijaylaxmi Trivedi.
\newblock A local {B}ertini theorem in mixed characteristic.
\newblock {\em Comm. Algebra}, 22(3):823--827, 1994.

\bibitem[Tri97]{TrivediLocalBertiniErratum}
Vijaylaxmi Trivedi.
\newblock Erratum: ``{A} local {B}ertini theorem in mixed characteristic''
  [{C}omm. {A}lgebra {\bf 22} (1994), no. 3, 823--827; {MR}1261007
  (94m:13002)].
\newblock {\em Comm. Algebra}, 25(5):1685--1686, 1997.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath}
\usepackage{verbatim}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage[dvipsnames]{xcolor}
\usepackage[all]{xy}
\usepackage{enumerate}
\usepackage[top=1in, bottom=1in, left=0.9in, right=0.9in]{geometry}
\usepackage{mathrsfs}

\usepackage{stmaryrd}
\usepackage{bbold}

\usepackage{soul}
\usepackage[colorlinks=true,citecolor=PineGreen,linkcolor=RawSienna]{hyperref}
\usepackage{mabliautoref}
\input{kmacros3.sty}

\numberwithin{equation}{section}

\usepackage{mabliautoref}

\renewcommand{\m}{\mathfrak{m}}

\newcommand{\pd}{\operatorname{pd}}
\newcommand{\fd}{\operatorname{fd}}

\renewcommand{\Spec}{\operatorname{Spec}}

\newcommand{\depth}{\operatorname{depth}}

\renewcommand{\leq}{\leqslant}
\renewcommand{\geq}{\geqslant}

\title{A Kunz-type characterization of regular rings via alterations}
\author{Linquan Ma}
\thanks{Ma was supported in part by NSF Grant DMS \#1836867/1600198.}
\author{Karl Schwede}
\thanks{Schwede was supported in party by NSF CAREER Grant DMS \#1252860/1501102 and NSF Grant \#1801849.}
\address{Department of Mathematics\\ Purdue University\\  West Lafayette\\ IN 47907}
\email{ma326@purdue.edu}
\address{Department of Mathematics\\ University of Utah\\ Salt Lake City\\ UT 84112}
\email{schwede@math.utah.edu}

\subjclass[2010]{14F18, 13A35, 14F17, 13D05, 13D45, 14B15, 14B05}
\keywords{multiplier ideals, projective dimension, regular rings, rational singularities}

\begin{document}

\begin{abstract}
We prove that a local domain $R$, essentially of finite type over a field, is regular if and only if for every regular alteration $\pi : X \to \Spec R$, we have that $\myR \pi_* \O_X$ has finite (equivalently zero in characteristic zero) projective dimension.
\end{abstract}

\maketitle

\section{Introduction}

In \cite{KunzCharacterizationsOfRegularLocalRings} Kunz proved that a Noetherian ring of characteristic $p > 0$ is regular if and only if the $e$-th iterated Frobenius map
\[
\xymatrix@R=3pt{
R \ar[r] & R \\
x \ar@{|->}[r] & x^{p^e}
}
\]
is flat for some, or equivalently every, $e > 0$. This is generalized in \cite{RodicioOnaResultAvramov}: the condition $\fd_RR^{(e)}<\infty$ (for some, or equivalently every, $e>0$) implies $R$ is regular, where $R^{(e)}$ denotes the target of the $e$-th Frobenius map. Moreover, the direct limit of $R^{(e)}$ is the perfection $R^\infty$ of $R$. Kunz's theorem can also be generalized using perfection: $R$ is regular if and only if $\fd_RR^{\infty}<\infty$ \cite{AberbachLiAsymptoticVanishingRegular,BhattIyengarMaRegularRingsPerfectoidAlgebras}.

However, in all these characterizations, the Frobenius map plays a prominent role and hence they do not extend to characteristic zero. In this paper, motivated by the connections between multiplier ideals and test ideals \cite{BlickleSchwedeTuckerTestAlterations}, we prove the following characterization of regularity using alterations.

\begin{mainthm*}
Suppose $(R, \fram)$ is a local domain essentially of finite type over a field.  Then the following conditions are equivalent.
\begin{itemize}
\item[(a)]  $R$ is regular.
\item[(b)]  For every regular alteration $\pi : X \to \Spec R$, $\pd_R \myR \pi_* \O_X < \infty$ (i.e., the derived image of the structure sheaf has finite projective dimension).
\end{itemize}
Moreover, when $R$ has characteristic $0$, the above are also equivalent to
\begin{itemize}
\item[(c)]  For every regular alteration $\pi : X \to \Spec R$, $\pd_R \myR \pi_* \O_X = 0$.
\end{itemize}
\end{mainthm*}

This result is also motivated by the fact that there is close connection between big Cohen-Macaulay algebras and $\myR \pi_* \O_X$ (recall that the latter is a Cohen-Macaulay complex, \cite{RobertsCohenMacaulayComplexes}).

In fact, this theorem in characteristic $p>0$ essentially follows from a result of Bhatt on killing cohomology using finite covers \cite{BhattDerivedsplintersinpositivecharacteristic} and the characterization of regularity using $R^+$, see \cite{AberbachLiAsymptoticVanishingRegular,BhattIyengarMaRegularRingsPerfectoidAlgebras}. Our main contribution is the characteristic zero case.  Meanwhile, it is quite natural to ask whether the same characterization of regularity holds in mixed characteristic:

\begin{question}
Suppose $(R, \fram)$ is an excellent local domain essentially of finite type over the integers. If for every regular alteration $\pi : X \to \Spec R$, we have that $\pd_R \myR \pi_* \O_X < \infty$, then is $R$ regular?
\end{question}

See \autoref{rmk.MixedCharacteristicSurface} for some additional discussion.  Finally, one should also compare with the classical result that if $R \subseteq S$ is a flat local extension and $S$ is regular, then so is $R$, see \cite[Theorem 23.7]{MatsumuraCommutativeRingTheory}.

\subsubsection*{Acknowledgements}  We thank Bhargav Bhatt and Srikanth Iyengar for comments on a previous draft of this article.

\section{Preliminaries}
Suppose $R$ is a Noetherian ring. For a bounded above cochain complex of $R$-modules $C$ ($\dots \to C^{-1} \to C^0 \to C^1 \to \dots$), we define the \emph{projective dimension of $C$}, denoted $\pd_R C$ to be
\[
\inf \{ \sup \{ i \;|\; P^{-i} \neq 0 \} \;|\; P^{\bullet} \text{ is a projective resolution of $C$} \}.
\]
Here a projective resolution is a cochain complex of projective modules quasi-isomorphic to $C$. If $(R, \fram)$ is additionally local, we also define
\[
\depth C = \inf \{ i \;|\; H^i_{\fram}(C) \neq 0 \}.
\]
These are natural extensions of the projective dimension and the depth of a finitely generated $R$-module.

We will need the following result of Foxby and Iyengar, which is a vast generalization of the classical Auslander-Buchsbaum formula.
\begin{theorem}[\cite{FoxbyIyengarDepthAmplitude}]
\label{thm.BuchEisComplex}
Let $(R,\m)$ be a local ring and let $M$ and $P$ be complexes of $R$-modules.  If $\pd_R P$ is finite and ${\mathrm{H}}(P)$ is nonzero and finitely generated, then
\[
\depth_R M = \depth_R(M \otimes^{\bf L}_R P) + \pd_R P.
\]
\end{theorem}

We can now prove our main result in positive characteristic.

\begin{proof}[Proof of Main Theorem in characteristic $p>0$]
First of all by \autoref{thm.BuchEisComplex}, $$\pd_R \myR\pi_*\O_X+\depth \myR\pi_*\O_X=\depth R.$$
Since $\depth \myR\pi_*\O_X\geq 0$ (as $\myR\pi_*\O_X$ lives in positive degree), we know that for all regular alteration, $\pd_R\myR\pi_*\O_X\leq \depth R$.

By \cite[Theorem 1.5]{BhattDerivedsplintersinpositivecharacteristic}, for every regular alteration $\pi$: $X\to \Spec R$, there exists another regular alteration $\pi'$: $Y\xrightarrow{f} X\to \Spec R$ such that the map $\tau_{\geq 1}\myR \pi_*\O_X\to \tau_{\geq 1}\myR \pi'_*\O_Y$, induced by the diagram of triangles below, is $0$.
\[\xymatrix{
\pi_*\O_X \ar[r] \ar[d] & \myR \pi_*\O_X \ar[r] \ar[d] & \tau_{\geq 1}\myR \pi_*\O_X \ar[d]^0 \ar[r]^{+1} & \\
\pi'_*\O_Y \ar[r] & \myR \pi'_*\O_Y \ar[r] & \tau_{\geq 1}\myR \pi'_*\O_Y \ar[r]_{+1} &
}
\]
Tensoring with $k=R/\m$ and taking cohomology, for all $i>\depth R$ we get:
\footnotesize
\[\xymatrix{
0=\myH^{-i-1}(\myR \pi_*\O_X\otimes^\myL k) \ar[r]& \myH^{-i-1} (\tau_{\geq 1}\myR \pi_*\O_X \otimes^\myL k)\ar[d]^0 \ar[r] & \myH^{-i}(\pi_*\O_X\otimes^\myL k) \ar[r] \ar[d] & 0=\myH^{-i}(\myR \pi_*\O_X\otimes^\myL k)  \\
0=\myH^{-i-1}(\myR \pi'_*\O_Y\otimes^\myL k) \ar[r]& \myH^{-i-1} (\tau_{\geq 1}\myR \pi_*\O_X \otimes^\myL k)   \ar[r] & \myH^{-i}(\pi'_*\O_Y\otimes^\myL k) \ar[r] & 0=\myH^{-i}(\myR \pi_*\O_X\otimes^\myL k)
}
\]
\normalsize

An easy diagram chasing shows that for all regular alteration $X$, we can find another regular alteration $Y$ such that $\myH^{-i}(\pi_*\O_X\otimes^\myL k)\to  \myH^{-i}(\pi'_*\O_Y\otimes^\myL k)$ is $0$.  By writing $R^+$ as the colimit over finite domain extensions of $R$, $R^+ = \lim_{S \supseteq R} S$, we see that $\myH^{-i}(R^+\otimes^\myL k)=\Tor_i^R(R^+,k)=0$ for all $i>\depth R$. Now by \cite[Corollary 3.5]{AberbachLiAsymptoticVanishingRegular} or \cite[Theorem 4.13]{BhattIyengarMaRegularRingsPerfectoidAlgebras}, $R$ is regular.
\end{proof}

We recall the following very useful result of Corso-Huneke-Katz-Vasconcelos.
\begin{theorem}\textnormal{(\cite[Corollary 3.3]{CorsoHunekeKatzVasconcelos})}
\label{thm.CorsoHunekeKatzVasconcelos}
Suppose that $(R, \fram)$ is a Noetherian local ring and that $I$ is integrally closed and $\fram$-primary. Then $M$ has
projective dimension less than $t$ if and only if $\Tor^R_t(R/I,M) = 0$.
\end{theorem}

We specialize it in the following corollary that we will use in the next section.
\begin{corollary}
\label{cor.CorsoHunekeKatzVasconcelos}
Suppose that $(R, \fram)$ is a Noetherian local ring and that $I$ is an integrally closed $\fram$-primary ideal of finite projective dimension.  Then $R$ is regular.
\end{corollary}
\begin{proof}
Since $I$ has finite projective dimension, we see that $\Tor^R_i(R/I,k) = 0$ for $i \gg 0$.  But now taking $M = k = R/\fram$ in the statement of \autoref{thm.CorsoHunekeKatzVasconcelos} we see that $k$ has finite projective dimension since $I$ is integrally closed and $\fram$-primary.  The result follows.
\end{proof}

\subsection{Multiplier ideals and multiplier submodules}

For references in this section, see \cite{LazarsfeldPositivity2,SchwedeTakagiRationalPairs,BlickleSchwedeTuckerTestAlterations}.

\begin{definition}[Multiplier submodules]
Suppose that $\pi : X \to \Spec R$ is a resolution of singularities.  Then the \emph{multiplier submodule} of $R$, denoted $\mJ(\omega_R)$ is just $\pi_* \omega_X \subseteq \omega_R$.  Here $\omega_R$ (respectively $\omega_X$) is the first nonzero cohomology of the dualizing complex.

We now generalize this a bit.  Suppose $R$ is a normal domain, $\Gamma \geq 0$ is a $\bQ$-Cartier divisor, and $\pi$ is a log resolution of $(X, \Gamma)$.  Then we define $\mJ(\omega_R, \Gamma) = \pi_* \O_X(\lceil K_X - \pi^* \Gamma \rceil)$.  If we choose $0 \neq f \in R$ and $t \in \bQ_{\geq 0}$, then we set $\mJ(\omega_R, \Gamma, f^t) = \mJ(\omega_R, \Gamma + t \Div(f))$.  Finally, if $\fra \subseteq R$ is an ideal and $\pi$ is a log resolution of $(R, \Gamma, \fra^t)$ with $\fra \cdot \O_X = \O_X(-G)$, then we define
\[
\mJ(\omega_R, \Gamma, \fra^t) = \pi_* \O_X(\lceil K_X - \pi^* \Gamma - t G \rceil) \subseteq \omega_R.
\]
All of this is independent of the choice of resolution.
\end{definition}

In the above, if $\Gamma$ is ever left out, it is treated as zero.

\begin{definition}[Multiplier ideals]
Suppose that $R$ is a normal domain, $\Delta \geq 0$ is a $\bQ$-divisor such that $K_R + \Delta$ is $\bQ$-Cartier, $\fra \subseteq R$ is an ideal and $t \in \bQ_{\geq 0}$, then we define the \emph{multiplier ideal}
\[
\mJ(R, \Delta, \fra^t) = \pi_* \O_X(\lceil K_X - \pi^*(K_R + \Delta) - tG\rceil)
\]
where $\pi : X \to \Spec R$ is a log resolution of $(R, \Delta, \fra^t)$ and $\fra \cdot \O_X  = \O_X(-G)$.
Again, this is independent of the choice of resolution.
\end{definition}

If $\Delta$ is left off, then it is treated as zero and if $\fra$ is left off, it is treated as $R$.

\section{The main result in characteristic zero}

We begin with the ``easy'' direction.

\begin{theorem}
Suppose that $(R, \fram)$ is a regular local ring essentially of finite type over a field of characteristic zero.  If $\pi : X \to \Spec R$ is a regular alteration, then $\pd_R \myR \pi_* \O_X = 0$.
\end{theorem}
\begin{proof}
Since $R$ is regular, the bounded complex  $\myR \pi_* \O_X$ has finite projective dimension. By \autoref{thm.BuchEisComplex}, taking $M = R$ and $P = \myR \pi_* \O_X$ we have that
\[
\pd_R \myR \pi_* \O_X + \depth_R(\myR \pi_* \O_X) = \dim R.
\]
By the Matlis-dual version of Grauert-Riemenschneider vanishing \cite{GRVanishing}, $H^i_{\fram}(\myR \pi_* \O_X) = 0$ for all $i < \dim R$ and hence $\depth_R(\myR \pi_* \O_X) \geq \dim R$. Thus $\pd_R \myR \pi_* \O_X\leq 0$. But clearly $\pd_R \myR \pi_* \O_X\geq 0$ since $\myH^0(\myR \pi_* \O_X)\neq 0$. The result follows.
\end{proof}

\begin{lemma}
\label{lem.FPDImpliesCM}
Suppose $(R, \fram)$ is a local domain essentially of finite type over a field of characteristic zero and that $\pi : X \to \Spec R$ is a resolution of singularities.  If $\pd_R \myR \pi_* \O_X < \infty$, then $R$ is Cohen-Macaulay.
\end{lemma}
\begin{proof}
Since $\pd_R \myR \pi_* \O_X < \infty$, we see that the injective dimension of the Grothendieck dual, $\myR \pi_* \omega_X^{\mydot} \cong \pi_* \omega_X[\dim R]$ (by Grauert-Riemenschneider vanishing), is finite.  But then $\pi_* \omega_X$ is a finitely generated $R$-module of finite injective dimension and so $R$ is Cohen-Macaulay by Bass' question \cite{PeskineSzpiroDimensionProjective,HochsterTopicsInTheHomologicalTheory,RobertsIntersectionTheorems}.
\end{proof}
\begin{proof}[Alternate proof]
By the Matlis-dual version of the Grauert-Riemenschneider vanishing, we see that $H^i_{\fram}(\myR \pi_* \O_X) = 0$ for all $i < \dim R$.  Hence $\depth \myR \pi_* \O_X = \dim R$. Note also that $\pd_R \myR \pi_* \O_X \geq 0$ since $\myH^0(\myR \pi_* \O_X)\neq 0$. Thus we have
\[
\depth R = \depth(\myR \pi_* \O_X) + \pd_R(\myR \pi_* \O_X) \geq \dim R
\]
by \autoref{thm.BuchEisComplex} and hence $R$ is Cohen-Macaulay.
\end{proof}

We are ready to prove the following characterization of rational singularities, this result is an important step towards proving the main theorem and is interesting in its own right.

\begin{theorem}
\label{thm.RationalSing}
Suppose $(R, \fram)$ is a local domain essentially of finite type over a field of characteristic zero.  Let $\pi : X \to \Spec R$ be a resolution of singularities.  Then $R$ has rational singularities if and only if $\pd_R \myR \pi_* \O_X<\infty$.
\end{theorem}
\begin{proof}
If $R$ has rational singularities then obviously $\pd_R \myR \pi_* \O_X<\infty$ since $\myR \pi_* \O_X\cong R$. We now assume that $\pd_R \myR \pi_* \O_X<\infty$. We already see that $R$ is Cohen-Macaulay by \autoref{lem.FPDImpliesCM}. Hence, it is sufficient to show that $\pi_* \omega_X = \omega_R$.

So we suppose $\pi_* \omega_X \neq \omega_R$. By choosing a minimal prime $P$ of $\Supp (\omega_R/\pi_*\omega_X)$ and replacing $R$ by $R_P$, we may assume $R$ has rational singularities on the punctured spectrum (i.e., $\omega_R/\pi_*\omega_X$ has finite length). Since $\pi_* \omega_X$ has finite injective dimension (see the proof of \autoref{lem.FPDImpliesCM}), by \cite[Theorem 2.9]{SharpFinitelyGeneratedmodulesFiniteInjectiveDimension}, $\Hom_R(\omega_R, \pi_* \omega_X)$ has finite projective dimension.  But
\[
\Hom_R(\omega_R, \pi_* \omega_X) = \pi_* \sHom_{\O_X}(\pi^* \omega_R, \omega_X).
\]
Now $\sHom_{\O_X}(\pi^* \omega_R, \omega_X)$ is a rank $1$ reflexive sheaf on $X$. Since $X$ is regular, $\sHom_{\O_X}(\pi^* \omega_R, \omega_X)$ is locally free and so its pushforward, which is isomorphic to $$\Hom_R(\omega_R, \pi_* \omega_X) \subseteq \Hom_R(\omega_R, \omega_R) \subseteq R,$$ is an integrally closed ideal. Since our assumption is $0\neq \omega_R/\pi_*\omega_X$ has finite length, it follows that $\Hom_R(\omega_R, \pi_* \omega_X)\neq R$ is an $\m$-primary integrally closed ideal. But then by \autoref{cor.CorsoHunekeKatzVasconcelos}, $\pd_R\Hom_R(\omega_R, \pi_* \omega_X)<\infty$ already implies $R$ is regular and thus $\pi_* \omega_X =\omega_R$ which is a contradiction.
\end{proof}
\begin{remark}
Bhargav Bhatt communicated to us an alternate proof of \autoref{thm.RationalSing}, which we now sketch.
Since $\myR \pi_* \O_X$ is a perfect complex, there exists a trace map
\[
\myR \Hom_R(\myR \pi_* \O_X, \myR\pi_* \O_X) \to R.
\]
On the other hand, we have the map $\myR \pi_* \O_X \to \Hom_R(\myR \pi_* \O_X, \myR\pi_* \O_X)$ coming from $\O_X$'s left multiplication action on itself.  We have the composition
\[
R \to \myR \pi_* \O_X \to \myR \Hom_R(\myR \pi_* \O_X, \myR\pi_* \O_X) \to R
\]
which is an isomorphism generically (on the open set where $\pi$ is an isomorphism), hence an isomorphism.  But then $R$ has rational singularities by \cite{KovacsRat} (note that that result still utilizes Grauert-Riemenschneider vanishing).
\end{remark}

\subsection{An aside on multiplier ideals}

We assume the following is essentially well known to experts, but we do not know a reference.

\begin{proposition}
\label{prop.MultiplierIdeal}
Suppose $(R,\fram)$ is a normal local domain essentially of finite type over a field of characteristic zero. Suppose $0\neq f\in R$ such that $\Div_R(f)$ is reduced. Fix $N\geq 0$ and let $S=R[f^{1\over N+1}]$ be the normal cyclic cover. Then $\mJ(\omega_S)$ has an $R$-summand isomorphic to $\mJ(\omega_R, f^{N \over N+1})$.
\end{proposition}

\begin{proof}
Since $\Div_R(f)$ is reduced, $S$ is regular in codimension $1$ and hence $S$ is normal. Choose $-K_R$ effective.
By \cite[Theorem 9.5.42]{LazarsfeldPositivity2} (see also \cite[Theorem 8.1]{BlickleSchwedeTuckerTestAlterations}) we see that
\[
\mJ(R, -K_R + {N \over N+1} \Div_R f) = R \cap \mJ(S, -\Ram_{S/R}  - \rho^* K_R + {N \over N+1} \Div_S f ).
\]
Again since $\Div_R f$ is reduced, we see that $\Ram_{S/R} = {N \over N+1} \Div_S f$ and hence
\[
\mJ(R, -K_R + {N \over N+1} \Div_R f) \subseteq \mJ(S, -\rho^* K_R).
\]
On the other hand, by \cite[Theorem 8.1]{BlickleSchwedeTuckerTestAlterations}, we have a splitting (up to scalars)
\[
\begin{array}{rl}
& \Tr(\mJ(S, -\rho^* K_R)) \\
= & \Tr(\mJ(S, -\rho^* K_R - \Ram_{S/R} + {N \over N+1} \Div_S f )) \\
= & \mJ(R, -K_R + {N \over N+1} \Div_R f)\\
 = & \mJ(\omega_R, f^{N \over N+1}).
\end{array}
\]
But we have
\[
\begin{array}{rl}
& \mJ(S, -\rho^* K_R)\\
 = & \mJ(S, -\rho^* K_R - \Ram_{S/R} + {N \over N+1} \Div_S f ) \\
 = & \mJ(S, -K_S + {N \over N+1} \Div_S f ) \\
 = & \mJ(\omega_S, f^{N \over N+1}).
 \end{array}
\]
We have just shown that $\mJ(\omega_S, f^{N \over N+1})$ has an $R$-summand isomorphic to $\mJ(\omega_R, f^{N \over N+1})$.  But even as an $S$-module $\mJ(\omega_S, f^{N \over N+1}) = f^{N \over N+1} \mJ(\omega_S) \cong \mJ(\omega_S)$, and hence $\mJ(\omega_S)$ has an $R$-summand isomorphic to $\mJ(\omega_R, f^{N \over N+1})$.
\end{proof}

\subsection{Proof of Main Theorem in characteristic zero}

We now complete the proof of our main result in characteristic zero.

\begin{theorem}
Suppose $(R, \fram)$ is a local domain essentially of finite type over a field of characteristic zero.  Suppose that for every regular alteration $\pi : X \to \Spec R$, $\pd_R \myR \pi_* \O_X<\infty$.  Then $R$ is regular.
\end{theorem}
\begin{proof}
By \autoref{thm.RationalSing}, we already know that $R$ has rational singularities.  Choose $N >0$ so that $\mJ(\omega_R, \fram^N)\neq \omega_R$.  Then choose a general $f \in \fram^{N+1}$ and by \cite[Proposotion 9.2.28]{LazarsfeldPositivity2} we know that $\mJ(R, -K_R + {N \over N+1} \Div_R f ) = \mJ(R, -K_R, \fram^N) = \mJ(\omega_R, \fram^N)$.

Consider the normal cyclic cover $S = R[f^{1\over N+1}]$. Since $f$ is general, $\Div_R(f)$ is reduced and by \autoref{prop.MultiplierIdeal}, we know that $\mJ(\omega_S)$ has an $R$-summand isomorphic to $\mJ(\omega_R, f^{N \over N+1})$.

Next consider a resolution of singularities $\pi$: $X\to \Spec S$, then the composition $X\to \Spec S\to \Spec R$ is a regular alteration. Moreover, $\pi_*\omega_X=\mJ(\omega_S)$ has finite injective dimension over $R$ (because $\pi_*\omega_X[\dim R]$ is the Grothendieck dual of $\myR\pi_*\O_X$), so does its direct summand $\mJ(\omega_R, f^{N \over N+1})$. Therefore by \cite[Theorem 2.9]{SharpFinitelyGeneratedmodulesFiniteInjectiveDimension}, $$\Hom_R(\omega_R, \mJ(\omega_R, f^{N \over N+1})) \subseteq \Hom_R(\omega_R, \omega_R) \cong R$$ has finite projective dimension.  Since $\mJ(\omega_R, f^{N \over N+1})=\mJ(\omega_R,\m^N)$ agrees with $\omega_R$ except at the origin (where it \emph{does not} agree).  Thus
$\Hom_R(\omega_R, \mJ(\omega_R, f^{N \over N+1}))$ lacks the identity map $\omega_R \to \omega_R$ and hence it is identified with an $\fram$-primary ideal of $R$.

Next we show that $\Hom_R(\omega_R, \mJ(\omega_R, f^{N \over N+1})) \subseteq R$ is an integrally closed ideal.  Take a log resolution of singularities $\pi : X \to \Spec R$ of $(R, \Div_R(f))$. By definition we have $\mJ(\omega_R, f^{N \over N+1}) = \pi_* \O_X(\lceil K_X - {N \over N+1} \Div_X(f)\rceil )$.  Thus
\[
\begin{array}{rl}
& \Hom_R(\omega_R, \mJ(\omega_R, f^{N \over N+1}))  \\
= &\Hom_R(\omega_R, \pi_* \O_X(\lceil K_X - {N \over N+1} \Div_X(f)\rceil )) \\
= & \pi_* \sHom_{\O_X}(\pi^* \omega_R, \O_X(\lceil K_X - {N \over N+1} \Div_X(f)\rceil )).
\end{array}\]
As in the proof of \autoref{thm.RationalSing}, since $\sL := \sHom_{\O_X}(\pi^* \omega_R, \O_X(\lceil K_X - {N \over N+1} \Div_X(f)\rceil ))$ is a rank 1 reflexive sheaf and $X$ is regular, $\sL$ is invertible.  Thus $\Hom_R(\omega_R, \pi_* \O_X(\lceil K_X - {N \over N+1} \Div_X(f)\rceil ))$ is an integrally closed $\m$-primary ideal of finite projective dimension.  Therefore $R$ is regular by
\autoref{cor.CorsoHunekeKatzVasconcelos}.
\end{proof}

\begin{remark}
\label{rmk.MixedCharacteristicSurface}
We believe that the above proof can be run (essentially without change) for excellent surfaces even in mixed characteristic.
The key facts we need are that Grauert-Riemenschneider still holds for excellent surfaces \cite[Corollary 2.10]{LipmanDesingularizationOf2Dimensional} and that we can choose a general element $f$ in $\fram^{N+1}$ so that $\mJ(R, f^{N/N+1})$ is $\fram$-primary \cite{TrivediLocalBertini,TrivediLocalBertiniErratum} (using that $R$ is regular outside of the origin since we may reduce to the case that $R$ is normal).
\end{remark}
\bibliographystyle{alpha}
\bibliography{MainBib}
\end{document}